OpenStudy (anonymous):

From y=1/2[x]+4/5 , what are the values of a, h, and k? A)a = 2, h = 0, k = 4/5 B)a = 2, h = 4, k = 5 C)a =1/2 , h = 0, k = 1 D)a = 1/2 , h = 0, k = 4/5 E)a = 1, h = 0, k = 0

OpenStudy (anonymous):

@Klip @iGreen someone .-.

OpenStudy (igreen):

Sorry..I don't know what it's asking..

OpenStudy (igreen):

The equation is in the slope-form of "y = mx + b".

OpenStudy (anonymous):

prob of braacket function..???

OpenStudy (anonymous):

@*Summer* bracket function problem..?

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

\[y=1/2(x)+4/5\] where 1/2 =a x =h 4/5 =k